# Mapping and Influencing the Political Ideology of Large Language Models using Synthetic Personas

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Language models
@@ 3. METHODOLOGY
We utilize PersonaHub, a collection of 1 billion diverse personas111with 200,000 publicly accessible through Huggingface, to examine how persona descriptions shape the political orientations of LLMs. To systematically evaluate these influences, we employ the PCT, which assesses political perspectives through 62 distinct statements across six key domains. While previous research by Lunardi et al. (2024) has shown that LLMs’ responses may vary based on question phrasing, we maintain the original PCT formulations to ensure methodological consistency. Our analysis includes 12.4 million categorical responses (200.000 personas x 62 statements), collected by prompting models to select from predefined political stance options.
@@
Our experimental setup comprises of two phases. First, we establish a baseline by having each model complete the PCT while impersonating each persona description. Second, we investigate ideological malleability by injecting explicit political descriptors (“right authoritarian” or “left libertarian”) to the persona descriptions and measuring the resulting shifts in political positioning.
@@ X Variable Y Variable
  • Note. $\Delta\mu$ = mean difference; $\sigma$ = standard deviation; WSR = Wilcoxon signed rank test; d = Cohen’s d effect size; CI = confidence interval.
  • p ¡ .05, *p ¡ .01, ***p ¡ .001.
@@
For the “left-libertarian” condition, the shifts were generally smaller in magnitude. Llama and Qwen showed minimal movement on the horizontal axis ($\Delta\mu_{x}=0.16$ and $-0.15$, respectively) and moderate changes on the vertical axis ($\Delta\mu_{y}$ = 0.68 and 0.47). This discrepancy might stem from the models’ pre-existing left-libertarian inclinations, making further adjustments less pronounced. This asymmetry may reflect biases in pretraining data, where left-libertarian narratives are more prevalent, leading to stronger internal representations of these ideologies. Additionally, the directional differences in responsiveness between left-right shifts versus authoritarian-libertarian shifts reveal that LLMs’ internal conceptualization of political ideologies may not fully align with the political compass framework.
@@ 5. CONCLUSION
In this study we investigated how adaptable LLMs are when impersonating certain profiles that may carry certain political ideologies. Our experimental results show that persona descriptions significantly influence LLMs’ political compass placements, with models exhibiting varying degrees of malleability to both implicit persona characteristics and explicit ideological prompting. This has important implications for understanding and controlling political bias in language models.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: XX.XXXX/XXXXXXX.XXXXXXXisbn: 979-8-4007-1331-6/25/04ccs: Information systems Language models
@@ 3. METHODOLOGY
We utilize PersonaHub, a collection of 1 billion diverse personas111with 200,000 publicly accessible through Huggingface, to examine how persona descriptions shape the political orientations of LLMs. To systematically evaluate these influences, we employ the PCT, which assesses political perspectives through 62 distinct statements across six key domains. While previous research by Lunardi et al. (2024) has shown that LLMs’ responses may vary based on question phrasing, we maintain the original PCT formulations to ensure methodological consistency. Our analysis includes 12.4 million categorical responses (200,000 personas x 62 statements), collected by prompting models to select from predefined political stance options.
@@
Our experimental setup comprises of two phases. First, we establish a baseline by having each model complete the PCT while impersonating each persona description. Second, we investigate ideological malleability by injecting explicit political descriptors (“right authoritarian” or “left libertarian”) to the persona descriptions and measuring the shifts in their political positioning. [^2] Code and data are openly available at https://github.com/d-lab/llm-political-personas
@@ X Variable Y Variable
  • Note. $\Delta\mu$ = mean difference; $\sigma$ = standard deviation; WSR = Wilcoxon signed rank test; d = Cohen’s d effect size; CI = confidence interval; ***p ¡ .001.
@@
For the “left-libertarian” condition, the shifts were generally smaller in magnitude. Llama and Qwen showed minimal movement on the horizontal axis ($\Delta\mu_{x}=0.16$ and $-0.15$, respectively) and moderate changes on the vertical axis ($\Delta\mu_{y}$ = 0.68 and 0.47). This discrepancy might stem from the models’ pre-existing left-libertarian inclinations, making further adjustments less pronounced. This asymmetry may reflect biases in pretraining data, where left-libertarian narratives are more prevalent, leading to stronger internal representations of these ideologies. Additionally, the differences in responsiveness between left-right shifts versus authoritarian-libertarian shifts reveal that LLMs’ internal conceptualization of political ideologies may not fully align with the political compass framework.
@@ 5. CONCLUSION
In this study we investigated how adaptable LLMs are when impersonating profiles that may carry certain political ideologies. Our results show that persona descriptions significantly influence LLMs’ political compass placements, with models exhibiting varying degrees of malleability to both implicit persona characteristics and explicit ideological prompting. This has important implications for understanding and controlling their political biases.
@@
Acknowledgments
This work is partially supported by the Australian Research Council (ARC) Training Centre for Information Resilience (Grant No. IC200100022) and by an ARC Future Fellowship Project (Grant No. FT240100022).